The length of a rectangle is 6 cm more than two times the width. If the perimeter of the rectangle is 36 cm, what are its dimensions?
Answer by checkley77(12844)

L=2W+6
2L+2W=36
2(2W+6)+2W=36
4W+12+2W=36
8W=36-12
6W=24
W=24/6
W=4 cm. ANS. FOR THE WIDTH.
L=2*4+6=8+6=14 cm. ANS.FOR THE LENGTH.
PROOF:
2*14+2*4=36
28+8=36
36=36